1- Find the domain of function f(X)=ln (-2/X2-x-6) +√x2 -1



2- Find the inverse function of the function f(X)=X2 -4x +5; X €(3,4)



3- construct the tangent line to the graph of the function f(X)= 4x.√x-2 .√x which is parallel to the line y=x.

1
Expert's answer
2022-05-26T07:16:47-0400

1.


"f(x)=\ln(-\dfrac{2}{x^2-x-6})+\sqrt{x^2-1}""x^2-x-6<0""x^2-1\ge0""(x+2)(x-3)<0""x\le -1\ or\ x\ge1"

Domain:"(-2, -1]\cup[1,3)"


2.



"𝑓(𝑥)=x^2−4x+5,x\in [3, 4]""x_v=-\dfrac{-4}{2(1)}=2"

The function "f" increases on "(3, 4)"



"f(3)=(3)^2\u22124(3)+5=2""f(4)=(4)^2\u22124(4)+5=5"

Domain: "[3, 4]"

Range: "[2, 5]"



"y=x^2-4x+5, 3\le x\le4"

Change "x" and "y"

"x=y^2-4y+5, 3\le y\le 4"

Solve for "y"

"y^2-4y+4=x-1""(y-2)^2=x-1"

Since "3\le y\le 4"



"y-2=\sqrt{x-1}"





"y=2+\sqrt{x-1}"

Then



"f^{-1}(x)=2+\sqrt{x-1}"


Domain: "[2, 5]"

Range: "[3, 4]"


3.



"f(x)=4x\sqrt{x}-2\sqrt{x}"

Domain: "[0, \infin)"



"f'(x)=4(\dfrac{3}{2})\sqrt{x}-\dfrac{2}{2\sqrt{x}}=\dfrac{6x-1}{\sqrt{x}}""slope=f'(x)=\dfrac{6x-1}{\sqrt{x}}=1""6x-\sqrt{x}-1=0""(3\sqrt{x}+1)(2\sqrt{x}-1)=0"

Since "\sqrt{x}\ge0," we take "2\sqrt{x}-1=0"



"\sqrt{x}=\dfrac{1}{2}""x=\dfrac{1}{4}""f(\dfrac{1}{4})=4(\dfrac{1}{4})\sqrt{\dfrac{1}{4}}-2\sqrt{\dfrac{1}{4}}=-\dfrac{1}{2}"

The tangent line to the graph is



"y+\dfrac{1}{2}=x-\dfrac{1}{4}""y=x-\dfrac{3}{4}"